In a solution 0.200 kg each of water and ethanol are present. Calculate the mole fraction of each
component.
1
Expert's answer
2015-05-22T10:26:27-0400
We have solution of water (200 g) and ethanol (200 g).

Let's count moles of water and athanol in solution:
Moles (water) = m(water) / Mr(water) = 200 / 18 = 11 mol.
Moles (ethanol) = m(ethanol) / Mr(ethanol) = 200 / 46 = 4 mol.

Mole fraction = moles(compound) / moles(total)
Mole fraction(water) = moles(water) / moles(total) = 11 / (11+4) = 0.73
Mole fraction(ethanol) = moles(ethanol) / moles(total) = 4 / (11+4) = 0.27